OpenStudy (anonymous):

find the det(A), where A is a square matrix satisfying the property that A(transpose)A= I

OpenStudy (anonymous):

\[A^{t}A=I\]

OpenStudy (zarkon):

for square matrices DET(AB)=Det(A)Det(B)

OpenStudy (zarkon):

I guess I should add that Det(I)=1 and Det(A)=Det(A^T)

OpenStudy (anonymous):

so i have \[\det(A ^{t} A)=\det(I)\] \[\det(A ^{t}) \det(A) =\det(I)\]

OpenStudy (zarkon):

\[\det(A ^{t}) \det(A) =\det(I)=1\] \[\det(A) \det(A) =1\] \[\det(A)^2=1\] \[\det(A)=\pm\sqrt{1}=\pm1\]